OpenStudy (anonymous):

For the sequence an=3n-5, what is the value of a30?

jimthompson5910 (jim_thompson5910):

Replace every n with 30, then evaluate \[\Large a_{n} = 3n - 5\] \[\Large a_{30} = 3*30 - 5\] \[\Large a_{30} = 90 - 5\] \[\Large a_{30} = 85\] So the 30th term of this sequence is 85